# Always-convex harmonic shears

### v1

@@ Abstract.
We determine completely the analytic functions $\varphi$ in the unit disk ${\mathbb{D}}$ such that for all (normalized) orientation-preserving harmonic mappings $f=h+\overline{g}$ which satisfy $h+g=\varphi$, the condition that $f$ maps ${\mathbb{D}}$ onto a convex domain holds. As a consequence, we obtain the following more general result: for a given complex number $\eta$, with $|\eta|=1$, we characterize those holomorphic mappings $\varphi$ in ${\mathbb{D}}$ such that every harmonic function $f=h+\overline{g}$ as above with $h-\eta g=\varphi$ maps ${\mathbb{D}}$ onto a convex domain.
@@
M. Wołoszkiewicz-Cyll was supported by a grant of National Science Center, Poland (NCN), 2022/06/X/ST1/01611.
July 2, 2025
@@ Introduction
A domain $\Omega$ in the complex plane is said to be convex in the direction $t\in[0,\pi)$ if the intersection of $\Omega$ with any line parallel to the line through 00 and $e^{it}$ is either an interval or empty. A convex domain is a domain convex in every direction. The locally univalent complex-valued function $f$ in the unit disk ${\mathbb{D}}$ is convex in the $t$ direction (resp. convex) if $f({\mathbb{D}})$ is convex in the direction $t$ (resp. convex). Since every domain convex in some direction is a close-to-convex domain, any locally univalent analytic function convex in some direction is univalent (see, for instance, [22] and the references therein).
@@
Given any holomorphic function $f$ in the unit disk with $f(0)=1-f^{\prime}(0)=0$ and such that $f({\mathbb{D}})$ is convex in the direction $t$, the rotation $f_{\gamma}(z)=e^{-\gamma i}f(e^{\gamma i}z)$, with $\gamma=t-\pi/2$, maps ${\mathbb{D}}$ onto a domain convex in the vertical direction (that is, in the $\pi/2$ direction) and satisfies the same normalizations. Hence, we can focus on this latter family of analytic mappings.
@@
The class of all analytic functions $f$ in ${\mathbb{D}}$ which are convex in the vertical direction and satisfy $f(0)=1-f^{\prime}(0)=0$ is denoted by $CV(i)$. It seems that Fejer [8, p. 61] was the first to call attention to this family, though Robertson [23, 24] was the first to make an intensive study of $CV(i)$. According to [10, p. 196], it was a great surprise to the mathematical community to learn that $f\in CV(i)$ does not imply that the dilation $f_{r}$, $10<r<$ 1, defined by $f_{r}(z)=f(rz)/r$ is also a member of this class [14]. An explicit example of a function that is convex in the vertical direction but whose restriction to the disk $|z|<r$ does not have this property for any radius $\sqrt{2}-1<r<1$ was constructed by Goodman and Saff [11]. They conjectured that the radius $\sqrt{2}-1$ is best possible, which was later proved by Ruscheweyh and Salinas [25].
@@
A complex-valued function $f$ is harmonic if both the real and imaginary parts of $f$ are harmonic. Every harmonic mapping in the unit disk has a canonical decomposition $f=h+\overline{g}$, where $h$ and $g$ are analytic functions in ${\mathbb{D}}$ with $g(0)=0$. The mapping $f$ is locally univalent if and only if its Jacobian $J_{f}=|h^{\prime}|^{2}-|g^{\prime}|^{2}$ does not vanish [17]. It is said to be orientation-preserving if $|h^{\prime}|>|g^{\prime}|$ in ${\mathbb{D}}$ or, equivalently, if $h$ is locally univalent and the dilatation $\omega=g^{\prime}/h^{\prime}$ satisfies the condition $\omega({\mathbb{D}})\subset{\mathbb{D}}$. Otherwise, the locally univalent function is orientation-reversing. It is obvious that $f$ is orientation-reversing if and only if $\overline{f}$ is orientation-preserving.
Let $S_{H}$ be the class which consists of those orientation-preserving univalent harmonic mappings $f=h+\overline{g}$ in the unit disk normalized by the conditions $h(0)=g(0)=0$ and $h^{\prime}(0)=1$. The class $S_{H}$ is normal: every sequence of functions in $S_{H}$ has a subsequence that converges locally uniformly in the unit disk; but it is not compact, as the limit function is not necessarily univalent. There is one further normalization which produces a compact normal family. Specifically, the class $S_{H}^{0}=\{f=h+\overline{g}\in S_{H}\colon g^{\prime}(0)=0\}$ is compact (see [7, Ch. 5]). The class $\mathcal{S}$ of analytic and univalent functions $\varphi$ in the unit disk normalized by the conditions $\varphi(0)=1-\varphi^{\prime}(0)=0$ is a subfamily of $S_{H}^{0}$. Notice also that the dilatation $\omega=g^{\prime}/h^{\prime}$ of any $f\in S_{H}^{0}$ is a holomorphic function in the unit disk that fixes the origin and satisfies $\omega({\mathbb{D}})\subset{\mathbb{D}}$. We will say that any holomorphic function in ${\mathbb{D}}$ satisfying these two properties is a Schwarz function.
It might seem that the requirement on the normalizations $h(0)=g(0)=1-h^{\prime}(0)=g^{\prime}(0)=0$ could be restrictive somehow. However, this is not the case: we can always get the corresponding normalization by using the (invertible) transformation
(1) $f=h+\overline{g}\longmapsto F=\frac{f-f(0)}{h^{\prime}(0)}$
followed by the composition $f_{0}=\tau\circ F$, where
(2) $\tau(z)=\frac{z-\overline{a}\overline{z}}{1-|a|^{2}}\,,\quad a=\frac{g^{\prime}(0)}{\overline{h^{\prime}(0)}}\,,$
which is invertible as well.
@@
The shear construction, introduced by Clunie and Sheil-Small [5], is an effective method for producing univalent harmonic mappings in ${\mathbb{D}}$ with a specified dilatation onto domains convex in one direction by “shearing” a univalent analytic mapping along parallel lines. It can be easily verified that the version presented here is equivalent to the one in [5, Thm. 5.3].
@@
gives rise to the harmonic mapping $f=h+\overline{g}\in S_{H}^{0}$ which maps ${\mathbb{D}}$ onto a domain convex in the vertical direction and has dilatation $\omega$. Note that the hypothesis about the local univalence of such function $f$ in Theorem A follows from the condition $\omega({\mathbb{D}})\subset{\mathbb{D}}$ and the fact that $\varphi$ is locally univalent.
@@
The class $C_{H}^{0}$ consists of orientation-preserving harmonic mappings $f=h+\overline{g}$ in $\mathbb{D}$ such that $h(0)=g(0)=1-h^{\prime}(0)=g^{\prime}(0)=0$ and which map the unit disk onto a convex domain. It is known that every orientation-preserving harmonic mapping $f$ which maps the unit disk onto a convex domain is univalent [5, Thm. 5.7]. Therefore, $C_{H}^{0}\subset S_{H}^{0}$. Once more, the requirement on the normalizations of the functions $h$ and $g$ is not that restrictive, since the transformations (1) and (2) preserve the convexity.
@@
It is worth noting that the condition characterizing convexness, as stated in the theorem, is not easy to verify in practice. Nevertheless, Hengartner and Schober in [13] managed to show that it is satisfied by any (normalized) orientation-preserving harmonic mapping $f=h+\overline{g}$ obtained from the solution to
$\left\{\begin{array}[]{cc}h(z)+g(z)=\frac{-i}{2}\log\frac{1+iz}{1-iz}\\ g^{\prime}(z)/h^{\prime}(z)=\omega(z)\end{array}\right.\,,\quad z\in{\mathbb{D}}\,,\quad h(0)=g(0)=0\,,$
@@
Boyd, Dorff, Nowak, Romney, and Wołoszkiewicz-Cyll [2], adapted the arguments in [13] to prove that if $f=h+\overline{g}$ is as above and
$\left\{\begin{array}[]{cc}h(z)+g(z)=\frac{z}{1-z}\\ g^{\prime}(z)/h^{\prime}(z)=\omega(z)\end{array}\right.\,,\quad z\in{\mathbb{D}}\,,\quad h(0)=g(0)=0\,,$
then, again, $f$ maps the unit disk onto a convex domain for every Schwarz function $\omega$.
The effectiveness of the shear construction in producing $C_{H}^{0}$ functions is, however, limited. A simple example that justifies this assertion is the following.
Example 1.
The harmonic mapping $f=h+\overline{g}\in S_{H}^{0}$ obtained from the solution to
$\left\{\begin{array}[]{cc}h(z)+g(z)=z\\ \omega(z)=-iz\end{array}\right.\,,\quad z\in{\mathbb{D}}\,,\quad h(0)=g(0)=0\,,$
maps the unit disk onto a domain that is not convex. Actually, $f$ satisfies that $f(z)=-iF(iz)$, $z\in{\mathbb{D}}$, where $F$ is the function in the example in [7, p. 38]. The function $F$ has the range shown in [7, Figure 3.1, p. 39], which is not convex. Therefore, $f({\mathbb{D}})$ is not a convex domain.
@@
It should be noted that there exists a connection between the parameter $\theta$ and the function $\varphi$ in (4) in order to obtain always-convex harmonic shears of the function $\varphi$ in the $\theta$ direction: even though the shears of $\mathcal{H}$ in the $\pi/2$ direction are always-convex, the shears of $\mathcal{H}$ in the 00 direction are not, as the following example shows.
@@
Let $f_{0}=h_{0}+\overline{g_{0}}\in S_{H}^{0}$ be defined, for $z$ in ${\mathbb{D}}$, by
$f_{0}(z)=\frac{1}{2}\,\frac{2z-z^{2}}{(1-z)^{2}}+\frac{1}{2}\,\overline{\frac{z^{2}}{(1-z)^{2}}}=\frac{z}{(1-z)^{2}}-i\,{\rm Im}\left(\frac{z}{1-z}\right)^{2}\,.$
It is not difficult to check that $f_{0}$ maps the unit disk onto the domain that contains the origin and is bounded by the parabola with focus at $z=-1/2$ and directrix ${\rm Re}\,z=0$. That is, $f_{0}({\mathbb{D}})$ is not convex.
The functions $h_{0}$ and $g_{0}$ in the canonical decomposition of $f_{0}$ satisfy
@@
where $I(z)=z$ for $z$ in the unit disk.
This last example shows that there is a Schwarz function ($\omega=I$) for which the horizontal shear of $\mathcal{H}$ with dilatation $\omega$ does not map the unit disk onto a convex domain. Hence, as announced before, the shears of $\mathcal{H}$ in the 00 direction are not always-convex.
@@ 1. The use of rotations
Our main goal is to characterize the pairs $(\eta,\varphi)$, where $\eta=e^{2i\theta}$, $0<2\theta\leq 2\pi$, has modulus one and $\varphi\in\mathcal{S}$ such that any harmonic mapping $f=h+\overline{g}$ obtained from the solution to the system
@@
Several examples of admissible pairs of the form $(-1,\varphi)$ have been given above. Other examples of admissible pairs $(\eta,\varphi)$ for $\eta$ not necessarily equal to $-1$ are provided by the next proposition. Here, for $\xi\in\partial{\mathbb{D}}$ and $f\in S_{H}^{0}$ we define the rotation
@@
Since $F$ is not convex, the function $F_{\xi}$ is not convex either. However, a straightforward calculation shows that ($H^{\xi},G^{\xi}$) solve (4) with datum $\omega$ defined in ${\mathbb{D}}$ by $\omega(z)=\xi^{2}\omega_{0}(\xi z)$. This gives the desired contradiction and shows that if $(\eta,\varphi)$ is admissible then $(\eta\xi^{2},\varphi_{\overline{\xi}})$ is also admissible. It is obvious that the reverse implication also holds since the transformation $(\eta,\varphi)\mapsto(\eta\xi^{2},\varphi_{\overline{\xi}})$ can be reverted. $\Box$
@@
In order to answer this question, it is worth noting that if a given harmonic mapping $f=h+\overline{g}\in S_{H}^{0}$ has the property that $h-g$ is not convex in the vertical direction, then the harmonic function $h-\overline{g}$ cannot be convex. An easy proof of this assertion follows from the fact that if $h-\overline{g}$ is convex, then $h+e^{2ti}g$ must be convex in the $t$ direction for each $t\in[0,\pi)$, according to Theorem B. This property does not hold for $t=\pi/2$ unless $h-g\in CV(i)$. This simple observation is the key tool to prove our next result.
@@
$-\xi z=\frac{G^{\prime}(z)}{H^{\prime}(z)}=\frac{G^{\prime}(z)}{h_{0}^{\prime}(\xi z)}=\frac{g_{0}^{\prime}(\xi z)}{h_{0}^{\prime}(\xi z)}\,\frac{G^{\prime}(z)}{g_{0}^{\prime}(\xi z)}=\xi z\,\frac{G^{\prime}(z)}{g_{0}^{\prime}(\xi z)}\,,\quad z\neq 0\,.$
@@
must be convex as well. Recall that $\xi=e^{i\theta}$ for some $\theta\in[0,\pi)$ and, by [5, Thm. 5.7], the harmonic mapping $h_{0}-\xi^{2}\overline{g}_{0}$ is convex if and only if for all $t\in[0,\pi]$ the analytic functions $h_{0}+e^{2it}\overline{\xi}^{2}\,g_{0}$ are convex in the $t$ direction. This means that, in particular, for $t=\theta$, the function $h_{0}+g_{0}$ must be convex in the direction $\theta$. But $h_{0}+g_{0}=k$, where $k$ is the Koebe function defined in the unit disk by $k(z)=z/(1-z)^{2}$. The domain $k({\mathbb{D}})$ is convex only when $\theta=0$ and hence $\xi=1$.
In cases where $\varphi=\mathcal{L}_{\lambda}$ for some $|\lambda|=1$ with a positive imaginary part, we can repeat the same argument used for the case $\varphi=\mathcal{H}$ above. The statement $\xi=1$ will be proved once we prove that for all such $\lambda$, the vertical shear $f=h+\overline{g}$ of $\mathcal{L}_{\lambda}$ with dilatation $-I$ satisfies that $h-g$ is convex only in the horizontal direction.
@@
The result is now evident from the Schwarz-Christoffel formula and the fact that for all $z\in{\mathbb{D}}$, $(h-g)^{\prime}(z)=\overline{(h-g)^{\prime}(\overline{z})}$. However, we have decided to include for the sake of completeness an alternative way of concluding that $h-g$ is convex only in the horizontal direction.
@@
which transforms the unit disk onto the entire complex plane minus the slit on the real axis that contains the interval $(-\infty,(1+{\rm Re}\lambda)/2]$. Note that $(1+{\rm Re}\lambda)/2$ is a positive real number and that the principal branch of the logarithm maps the whole complex plane minus the negative part of the real axis conformally onto the strip $\{z\in{\mathbb{C}}:\,|{\rm Im}z|<\pi\}$ and the positive part of the real axis to the real axis. Hence, the mapping $h-g$ transforms the unit disk onto a strip with a horizontal slit from $\log((1+{\rm Re}\lambda)/2)$ to +$\infty$+ ∞ and is therefore convex only in the horizontal direction. $\Box$

### v2

@@ Abstract.
We determine completely the analytic functions $\varphi$ in the unit disk $\mathbb{D}$ such that for all (normalized) orientation-preserving harmonic mappings $f=h+\overline{g}$ produced by the shear construction with $h+g=\varphi$, the condition that each $f$ maps $\mathbb{D}$ onto a convex domain holds. As a consequence, we obtain the following more general result: for a given complex number $\eta$, with $|\eta|=1$, we characterize those holomorphic mappings $\varphi$ in $\mathbb{D}$ such that every harmonic function $f=h+\overline{g}$ as above with $h-\eta g=\varphi$ maps $\mathbb{D}$ onto a convex domain. The resulting functions are mappings onto a half-plane and mappings onto a strip, and the shear direction, determined by the parameter $\eta$ above, is parallel to the linear boundaries of the half-planes and strips.
@@
R. Hernández and M. J. Martín were partially supported by Spanish MICIU research projectPID2024-160185NB-I00 MICIU research project PID2024-160185NB-I00.; M. Wołoszkiewicz-Cyll was supported by a grant of National Science Center, Poland (NCN), 2022/06/X/ST1/01611.
@@ Introduction
A domain $\Omega$ in the complex plane is said to be convex in the $t$ direction, $t\in[0,\pi)$, if the intersection of $\Omega$ with any line parallel to the line through 0 and $e^{it}$ is either an interval or empty. A convex domain is a domain convex in every direction. The univalent complex-valued function $f$ in the unit disk $\mathbb{D}$} is convex in the $t$ direction (resp. convex) if $f({\mathbb{D}})$ is convex in the $t$ direction (resp. convex). Every domain convex in some direction is a close-to-convex domain (see, for instance, [24] and the references therein).
@@
Given any holomorphic function $f$ in the unit disk with $f(0)=1-f^{\prime}(0)=0$ and such that $f({\mathbb{D}})$ is convex in the direction $t$, the rotation $f_{\gamma}(z)=e^{-\gamma i}f(e^{\gamma i}z)$, with $\gamma=t-\pi/2$, maps $\mathbb{D}$} onto a domain convex in the vertical direction (that is, in the $\pi/2$ direction) and satisfies the same normalizations. Hence, we can focus on this latter family of analytic mappings, denoted by $CV(i)$.
@@
Fejer [9, p. 61] was the first to call attention to this family, though Robertson [27, 28] was the first to make an intensive study of $CV(i)$. According to [11, p. 196], it was a great surprise to the mathematical community to learn that $f\in CV(i)$ does not imply that the dilation $f_{r}$, $0<r<1$, defined by $f_{r}(z)=f(rz)/r$, is also a member of this class [15]. An explicit example of a function that is convex in the vertical direction but whose restriction to the disk $|z|<r$ does not have this property for any radius $\sqrt{2}-1<r<1$ was constructed by Goodman and Saff [12]. They conjectured that the radius $\sqrt{2}-1$ is best possible, which was later proved by Ruscheweyh and Salinas [29].
@@
A complex-valued function $f$ is harmonic if both the real and imaginary parts of $f$ are harmonic. Every harmonic mapping $f$ in the unit disk has a canonical decomposition $f=h+\overline{g}$, where $h$ and $g$ are analytic in $\mathbb{D}$ with $g(0)=0$. A locally univalent function $f$ is called orientation-preserving if its Jacobian $J_{f}=|h^{\prime}|^{2}-|g^{\prime}|^{2}$ is positive in $\mathbb{D}$ [18], so that $\omega=g^{\prime}/h^{\prime}$ is analytic in $\mathbb{D}$ and $\omega(\mathbb{D})\subset\mathbb{D}$. If $J_{f}<0$ in $\mathbb{D}$, $f$ is orientation-reversing. It is obvious that $f$ is orientation-reversing if and only if $\overline{f}$ is orientation-preserving.
Let $S_{H}$ be the class which consists of those orientation-preserving univalent harmonic mappings $f=h+\overline{g}$ in the unit disk normalized by the conditions $h(0)=g(0)=0$ and $h^{\prime}(0)=1$. The class $S_{H}$ is known to be normal whereas the class $S_{H}^{0}=\{f=h+\overline{g}\in S_{H}\colon g^{\prime}(0)=0\}$ is compact (see [8, Ch. 5]). The class $\mathcal{S}$ of analytic and univalent functions $\varphi$ in the unit disk normalized by the conditions $\varphi(0)=1-\varphi^{\prime}(0)=0$ is a subfamily of $S_{H}^{0}$. Notice that the requirement on the normalizations $h(0)=g(0)=1-h^{\prime}(0)=g^{\prime}(0)=0$ can be always obtained by applying appropriate (invertible) affine harmonic transformations as described in [8, Ch. 5].
The dilatation $\omega=g^{\prime}/h^{\prime}$ of any $f\in S_{H}^{0}$ is a holomorphic function in the unit disk that fixes the origin and satisfies $\omega(\mathbb{D})\subset\mathbb{D}$. We will say that any holomorphic function $\omega$ in $\mathbb{D}$ satisfying these two properties is a Schwarz function.
@@
The shear construction, introduced by Clunie and Sheil-Small [6], is an effective method for producing univalent harmonic mappings in $\mathbb{D}$} with a specified dilatation onto domains convex in one direction by “shearing” a univalent analytic mapping along parallel lines. It can be easily verified that the version presented here is equivalent to the one in [6, Thm. 5.3].
@@
gives rise to the harmonic mapping $f=h+\overline{g}\in S_{H}^{0}$ which maps $\mathbb{D}$} onto a domain convex in the vertical direction and has dilatation $\omega$.
@@
It is easy to check that, in terms of $\varphi$ and $\omega$, the function $f=h+\overline{g}$ obtained from (1) has the form
$f(z)={\rm Re\,}\varphi(z)+i\,{\rm Im\,}\int_{0}^{z}\varphi^{\prime}(\zeta)\frac{1-\omega(\zeta)}{1+\omega(\zeta)}\,d\zeta\,.$
It is known that every orientation-preserving harmonic mapping $f$ which maps the unit disk onto a convex domain is univalent [6, Thm. 5.7]. The class $C_{H}^{0}$ consists of harmonic mappings $f=h+\overline{g}\in S_{H}^{0}$ which map $\mathbb{D}$} onto a convex domain.
@@
It is worth noting that the condition characterizing convexness, as stated in the theorem, is not easy to verify in practice. Nevertheless, Hengartner and Schober in [14] managed to show that it is satisfied by any (normalized) orientation-preserving harmonic mapping $f=h+\overline{g}$ obtained from the solution to (1) with
$\varphi(z)=\frac{-i}{2}\log\frac{1+iz}{1-iz}\,,\quad z\in{\mathbb{D}}\,$,
@@
Boyd et al. [3], adapted the arguments in [14] to prove that if $f=h+\overline{g}$ is obtained from (1) with $\varphi(z)=z/(1-z)$ for $z\in{\mathbb{D}}$, then, again, $f$ maps the unit disk onto a convex domain for every Schwarz function $\omega$.
The effectiveness of the shear construction in producing $C_{H}^{0}$ functions is, however, limited. A simple example that justifies this assertion is provided by the choices $\varphi(z)=z$ and $\omega(z)=-iz$ in (1). The resulting function $f$ satisfies that f(z)=iF(iz)f(z)=-iF(iz), $z\in{\mathbb{D}}$, where $F$ is the function in the example in [8, p. 38]. The function $F$ has the range shown in [8, Figure 3.1, p. 39], which is not convex. Therefore, $f({\mathbb{D}})$ is not a convex domain.
@@
It should be noted that there exists a connection between the parameter $\theta$ and the function $\varphi$ in (2) in order to obtain always-convex harmonic shears of the function $\varphi$ in the $\theta$ direction: even though the shears of $\mathcal{H}$ in the $\pi/2$ direction are always-convex, the shears of $\mathcal{H}$ in the 0 direction are not, as the following fundamental example shows. We refer the reader to [26] for the details on the relevance on the function we now present. It is worth mentioning also the nice articles [19] and [25], where the authors obtain different important consequences by using the strengthness of Theorem A (among other results from the literature combined in an appropriate way).
@@
Let $f_{0}=h_{0}+\overline{g_{0}}\in S_{H}^{0}$ be defined, for $z$ in $\mathbb{D}$}, by the solution to
@@
where $I(z)=z$ for $z$ in the unit disk and $\mathcal{H}$ is as in (3).
In [26], it is proved that $f_{0}$ is not starlike and hence, it is not convex (see [26, Fig. 7] for a representation of the image of $f_{0}$, which is denoted by $f_{3}$ in the article mentioned).
This last example shows that there is a Schwarz function ($\omega=I$) for which the horizontal shear of $\mathcal{H}$ with dilatation $\omega$ does not map the unit disk onto a convex domain. Hence, as announced before, the shears of $\mathcal{H}$ in the 0 direction are not always-convex.
@@ 1. The use of rotations
As mentioned above, our main goal is to characterize the pairs $(\eta,\varphi)$, where $\eta=e^{2i\theta}$, $0<2\theta\leq 2\pi$, has modulus one and $\varphi\in\mathcal{S}$ such that any harmonic mapping $f=h+\overline{g}$ obtained from the solution to the system
@@
Several examples of admissible pairs of the form $(-1,\varphi)$ have been given above. Other examples of admissible pairs $(\eta,\varphi)$ for $\eta$ not necessarily equal to 1-1 are provided by the next proposition. Here, for $\xi\in\partial{\mathbb{D}}$ and $f\in S_{H}^{0}$ we define the rotation
@@
Since $F$ is not convex, the function $F_{\xi}$ is not convex either. However, a straightforward calculation shows that ($H^{\xi},G^{\xi}$) solve (2) with datum $\omega$ defined in $\mathbb{D}$} by $\omega(z)=\xi^{2}\omega_{0}(\xi z)$. This gives the desired contradiction and shows that if $(\eta,\varphi)$ is admissible then $(\eta\xi^{2},\varphi_{\overline{\xi}})$ is also admissible. It is obvious that the reverse implication also holds since the transformation $(\eta,\varphi)\mapsto(\eta\xi^{2},\varphi_{\overline{\xi}})$ can be inverted. $\Box$
@@
In order to answer this question, it is worth noting that if a given harmonic mapping $f=h+\overline{g}\in S_{H}^{0}$ has the property that hgh-g is not convex in the vertical direction, then the harmonic function $h-\overline{g}$ cannot be convex. An easy proof of this assertion follows from the fact that if $h-\overline{g}$ is convex, then $h+e^{2ti}g$ must be convex in the $t$ direction for each $t\in[0,\pi)$, according to Theorem B. This property does not hold for $t=\pi/2$ unless $h-g\in CV(i)$. This simple observation is the key tool to prove our next result.
@@
$\xi z=\frac{G^{\prime}(z)}{H^{\prime}(z)}=\frac{G^{\prime}(z)}{h_{0}^{\prime}(\xi z)}=\frac{g_{0}^{\prime}(\xi z)}{h_{0}^{\prime}(\xi z)}\,\frac{G^{\prime}(z)}{g_{0}^{\prime}(\xi z)}=\xi z\,\frac{G^{\prime}(z)}{g_{0}^{\prime}(\xi z)}\,,\quad z\neq 0\,.$
@@
must be convex as well. Recall that $\xi=e^{i\theta}$ for some $\theta\in[0,\pi)$ and, by [6, Thm. 5.7], the harmonic mapping $h_{0}-\xi^{2}\overline{g}_{0}$ is convex if and only if for all $t\in[0,\pi]$ the analytic functions $h_{0}+e^{2it}\overline{\xi}^{2}\,g_{0}$ are convex in the $t$ direction. This means that, in particular, for $t=\theta$, the function $h_{0}+g_{0}$ must be convex in the direction $\theta$. But $h_{0}+g_{0}=k$, where $k$ is the Koebe function defined in the unit disk by $k(z)=z/(1-z)^{2}$. The domain $k({\mathbb{D}})$ is convex only when $\theta=0$ and hence $\xi=1$.
In cases where $\varphi=\mathcal{L}_{\lambda}$ for some $|\lambda|=1$ with a positive imaginary part, we can repeat the same argument used for the case $\varphi=\mathcal{H}$ above. The statement $\xi=1$ will be proved once we prove that for all such $\lambda$, the vertical shear $f=h+\overline{g}$ of $\mathcal{L}_{\lambda}$ with dilatation I-I satisfies that hgh-g is convex only in the horizontal direction.
@@
The result is now evident from the Schwarz-Christoffel formula and the fact that for all $z\in{\mathbb{D}}$, $(h-g)^{\prime}(z)=\overline{(h-g)^{\prime}(\overline{z})}$. However, we have decided to include for the sake of completeness an alternative way of concluding that hgh-g is convex only in the horizontal direction.
@@
which transforms the unit disk onto the entire complex plane minus the slit on the real axis that contains the interval $(-\infty,(1+{\rm Re}\lambda)/2]$. Note that $(1+{\rm Re}\lambda)/2$ is a positive real number and that the principal branch of the logarithm maps the whole complex plane minus the negative part of the real axis conformally onto the strip $\{z\in{\mathbb{C}}:\,|{\rm Im}z|<\pi\}$ and the positive part of the real axis to the real axis. Hence, the mapping hgh-g transforms the unit disk onto a strip with a horizontal slit from -$\infty$ to $\log((1+{\rm Re}\lambda)/2)$ and is therefore convex only in the horizontal direction. $\Box$